OpenStudy (anonymous):

a rectangle porch has a perimeter of 49 ft and a width of 6.5 ft find the area of the porch in square yards

jagr2713 (jagr2713):

Length = (49 - 2*6.5)/2 = 36/2 = 18. Area = length * width = 18 * 6.5 = ?